OpenStudy (anonymous):

Iron-59 has a half-life of 45.1 days. How old is an iron nail if the Fe-59 content is 25% that of a new sample of iron?

OpenStudy (anonymous):

25=100 e^(-.693 t/45.1) Ln(25)=ln(100) -.693 t/45.1 ln(25)-ln(100)=-.693 t/45.1 ln(25/100)=-.693t/45.1 t= - 45.1/.693 ( ln .25) t=90.2 days I put and it was wrong

OpenStudy (anonymous):

@TacianaCampbelll Let me try... Ln(25)=ln(100) -.693 t/45.1 ln(25)-ln(100)=-.693 t/45.1 ln(25/100)=-.693t/45.1 t= - 45.1/.693 ( ln .25) t=90.2 days. \(\checkmark\) Is Correct !!

OpenStudy (arihangdu):

How this happen!!!!